OpenStudy (anonymous):

How do I find the inverse of a given function and check it by using 3 points? The function is f(x)= 4x+3

OpenStudy (anonymous):

So far I have as the inverse (x-3)/4 I'm just not sure how to check it with the 3 points Please tell me if I am correct.

OpenStudy (anonymous):

this should help

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

Thanks, Now do I just pick any #'s for X & Y?

OpenStudy (anonymous):

\[f(x)=4x+3\\ f^{-1}(x)=\frac{x-3}{4}\]

OpenStudy (anonymous):

i think what they want you to do is this say for example \[f(2)=4\times 2+3=11\] and \[f^{-1}(11)=\frac{11-3}{4}=2\]

OpenStudy (anonymous):

so since \[f)(2)=11,f^{-1}(11)=2\] then it is a good indication it is right repeats with 2 other numbers

OpenStudy (anonymous):

That's what I was thinking, I could use different #'s for each and just switch my input (x) for my output (y) for the inverse

OpenStudy (anonymous):

Thanks (: